# J. A. Aguilar-Saavedra Instituto de Física Teórica, IFT-UAM/CSIC, c/ Nicolás Cabrera 13-15, 28049 Madrid

### v1

@@ Abstract
The decay $H\to ZZ\to 4\ell$, with $\ell=e,\mu$, can be used to test quantum interference in analogy to the famous double-slit experiment. The observations corresponding to ‘covering a slit’ can be extracted from data in the $ee\mu\mu$ channel, while the observations when both ‘slits’ are open, which include the interference, are measured in the $4e/4\mu$ channels. Furthermore, this process offers a unique opportunity to investigate identical-particle effects at high energies, and provide the first evidence of identical-particle behaviour for muons. Sensitivity at the $4\sigma$ level could be achieved at the high-luminosity upgrade of the Large Hadron Collider.
@@
In $4e/4\mu$ final states with identical particles there are two possible pairings between opposite-sign leptons. (Each pairing corresponds to one of the two tree-level $H\to ZZ\to 4\ell$ Feynman diagrams, related by exchange of identical particles.) Let us consider one such pairing and label with a common subindex the leptons in the same pair, $(\ell_{1}^{+},\ell_{1}^{-})$, $(\ell_{2}^{+},\ell_{2}^{-})$. One can then define the two $Z$ bosons by their momenta, $p_{Z_{1}}=p_{\ell_{1}^{+}}+p_{\ell_{1}^{-}}$, $p_{Z_{2}}=p_{\ell_{2}^{+}}+p_{\ell_{2}^{-}}$. And, as before, label as $Z_{1}$ the one with largest invariant mass. The reference system is then constructed as described for distinguishable particles, and likewise the four-dimensional decay distribution has the form (1). However, the numerical values of the coefficients $a_{LM}^{1,2}$ and $c_{L_{1}M_{1}L_{2}M_{2}}$ differ from the $ee\mu\mu$ final state. In particular, $c_{111-1}$ and $c_{1010}$ are up to 50 times larger, depending on the phase space region. We will precisely use the value of these coefficients to demonstrate the effects of interference and identical particles.
Probing quantum interference. In the $ee\mu\mu$ final state there is a ‘correct’ pairing of opposite-sign leptons (by flavour), let us label it as ‘A’. The ‘wrong’ pairing is labelled as ‘B’. We can regard each pairing as a ‘slit’, and measure $c_{111-1}$ and $c_{1010}$ in each case. That would be the analogous to covering any of the two slits in the classical experiment. In addition, we can also perform the measurement counting each event twice, one for each pairing. The result obtained is the average of both pairings, which do not interfere.
@@
In the $4e/4\mu$ final states we can also count each event twice, one for each possible pairing. But the results obtained for $c_{111-1}$ and $c_{1010}$ are not the average of pairings A and B, as in the $ee\mu\mu$ channel. In other words, the observation for $4e/4\mu$ is not the sum of the observations for either of the two ‘slits’ in $ee\mu\mu$. This is a clear indication of the interference. In terms of Feynman diagrams, each pairing — or ‘slit’ — can be identified with a diagram, but note that our definitions and method to verify the interference do not rely on an expansion of the amplitude in terms of Feynman diagrams.
@@
This can be illustrated with a Monte Carlo calculation at the parton level. We generate $H\to ZZ\to 4\ell$ at the leading order with MadGraph [10]. For $ee\mu\mu$ we obtain
@@
Future prospects. We estimate the statistical sensitivity of these measurements at the high-luminosity upgrade of the LHC (HL-LHC), with a luminosity of 3 ab-1. We consider $gg\to H\to ZZ\to 4\ell$ at 14 TeV, normalising the Higgs cross section to 54.67 pb and using branching ratios to $ee\mu\mu$ and $4e/4\mu$ of $5.9\times 10^{-5}$ and $6.5\times 10^{-5}$, respectively [11]. Hadronisation and parton showering are performed with Pythia [12] and detector simulation with Delphes [13] using the default card for the HL-LHC. Electrons are selected in the pseudorapidity region $|\eta|\leq 2.5$, and with transverse momentum $p_{T}\geq 7$ GeV. Muons are selected with $|\eta|\leq 2.7$ and $p_{T}\geq 5$ GeV. We require two opposite-sign same-flavour leptons. The three higher-$p_{T}$ leptons must satisfy minimum thresholds of 20, 15 and 10 GeV, respectively. These event selection requirements are quite close to those in Ref. [14]. In addition, we select a window $120\leq m_{4\ell}\leq 130$ GeV for the four-lepton invariant mass. In this region, the background has a size around 1/4 of the signal [14, 15]. The expected number of events in the $ee\mu\mu$, $4e$ and $4\mu$ final states are 1144, 379 and 1106, respectively. At the reconstructed level, $c_{111-1}$ and $c_{1010}$ exhibit moderate deviations from the parton-level values. We collect them in Table 1. For completeness we also give the values for the electroweak four-lepton background, calculated at the tree level, in the mass window $120\leq m_{4\ell}\leq 130$ GeV.
@@
A final comment on the electroweak four-lepton background is in order. In the region of interest it is quite small, as mentioned; the coefficients are also quite similar in $ee\mu\mu$ and $4e/4\mu$ final states, as seen in Table 1. And they are of similar magnitude as the ones for $H\to ZZ$. Therefore, a background subtraction, which is out of the scope of the present work, seems feasible, and will only result in a mild increase of the statistical uncertainty in $c_{111-1}$ and $c_{1010}$, and a corresponding decrease of the significance with respect to the results presented here. Next-to-leading order corrections for the $ee\mu\mu$ final state are known [18], and provide a small shift in $c_{111-1}$ and $c_{1010}$ which is well below the statistical uncertainty.

### v2

@@ Abstract
The decay $H\to ZZ\to 4\ell$, with $\ell=e,\mu$, can be used to test quantum interference in analogy to the famous double-slit experiment. The observations corresponding to ‘covering a slit’ can be extracted from data in the $ee\mu\mu$ channel, while the observations when both ‘slits’ are open, which include the interference, are measured in the $4e/4\mu$ channels. In this way, this process offers a unique opportunity to investigate identical-particle effects at high energies, and provide the first evidence of identical-particle behaviour for muons. Sensitivity at the $4\sigma$ level could be achieved at the high-luminosity upgrade of the Large Hadron Collider.
@@
In $4e/4\mu$ final states with identical particles there are two possible pairings between opposite-sign leptons. (Each pairing corresponds to one of the two tree-level $H\to ZZ\to 4\ell$ Feynman diagrams, related by exchange of identical particles.) Let us consider one such pairing and label with a common subindex the leptons in the same pair, $(\ell_{1}^{+},\ell_{1}^{-})$, $(\ell_{2}^{+},\ell_{2}^{-})$. One can then define the two $Z$ bosons by their momenta, $p_{Z_{1}}=p_{\ell_{1}^{+}}+p_{\ell_{1}^{-}}$, $p_{Z_{2}}=p_{\ell_{2}^{+}}+p_{\ell_{2}^{-}}$. And, as before, label as $Z_{1}$ the one with largest invariant mass. The reference system is then constructed as described for distinguishable particles, and likewise the four-dimensional decay distribution has the form (1). [^1] In principle, higher-rank spherical harmonics may enter the angular distribution for $4e/4\mu$ final states. We have checked that the coefficients for $L=3$ are compatible with zero within Monte Carlo uncertainty (for several million events). This can be understood since a fermion pair in s-wave has total angular momentum $J=0,1$, and behaves like a scalar or spin-1 particle. However, the numerical values of the coefficients $a_{LM}^{1,2}$ and $c_{L_{1}M_{1}L_{2}M_{2}}$ differ from the $ee\mu\mu$ final state. In particular, $c_{111-1}$ and $c_{1010}$ are up to 50 times larger, depending on the phase space region. We will precisely use the value of these coefficients to demonstrate the effects of interference and identical particles.
Probing quantum interference. In the $ee\mu\mu$ final state there is a ‘correct’ pairing of opposite-sign leptons (by flavour), let us label it as ‘A’. But one can also consider a ‘wrong’ pairing, labelled as ‘B’, in which we define $p_{X_{1}}=p_{e^{+}}+p_{\mu^{-}}$, $p_{X_{2}}=p_{\mu^{+}}+p_{e^{-}}$. Neither $X_{1}$ nor $X_{2}$ correspond to any intermediate $Z$ boson in this final state, but $p_{X_{i}}$ and $p_{X_{2}}$ are time-like momenta, so we can boost the $\mu^{-}$ and $e^{-}$ momenta to the $X_{1}$ and $X_{2}$ rest frames, respectively, define the angles $\Omega_{1,2}$ and from the distribution (1) calculate coefficients $c_{111-1}$ and $c_{1010}$ — which admittedly do not have any interpretation in terms of spin. (The motivation to consider this ‘wrong’ pairing is the comparison with the $4e/4\mu$ final state, as we see below.) Furthermore, we can perform the measurement of $c_{111-1}$ and $c_{1010}$ counting each event twice, one for each pairing. In this case, the result obtained is the average of the results obtained for each pairing, i.e. an incoherent sum.
@@
In $4e/4\mu$ final states there is no ‘correct’ pairing but we can still count each event twice, one for each pairing. And the results obtained for $c_{111-1}$ and $c_{1010}$ are not the average of pairings A and B as obtained for the $ee\mu\mu$ channel. In other words, the observation for $4e/4\mu$ is not the incoherent sum of the observations for the two choices in $ee\mu\mu$. (That would be the case in the absence of interference, as it can be seen by symmetry arguments.) In terms of Feynman diagrams, each pairing in this final state corresponds to a diagram; however, our definitions and method to verify the interference by comparison with the $ee\mu\mu$ final state do not depend on expanding the amplitude in terms of diagrams.
@@
This discussion can be illustrated with a Monte Carlo calculation at the parton level in full phase space. We generate $H\to ZZ\to 4\ell$ at the leading order with MadGraph [10]. For $ee\mu\mu$ we obtain
@@
The analogy with the double-slit experiment follows since in $4e/4\mu$ final states each lepton pairing is a possibility (e.g. to reconstruct the virtual $Z$ bosons), which can be regarded as a ‘slit’. Physical results are the coherent sum of both, as it is evident from the amplitude expansion in Feynman diagrams. In order to make the interference effect manifest we have considered the coefficients $c_{111-1}$, $c_{1010}$ obtained when summing over both pairings, and compared with the same quantity for the $ee\mu\mu$ final state. Following the analogy, the result of ‘covering a slit’, that is, the result obtained with only one pairing, can be experimentally measured in $ee\mu\mu$. And it is found that in $4e/4\mu$ the result is not the average of both, again in analogy with the double-slit experiment.
@@
Future prospects. We estimate the statistical sensitivity of these measurements at the high-luminosity upgrade of the LHC (HL-LHC), with a luminosity of 3 ab-1. We consider $gg\to H\to ZZ\to 4\ell$ at 14 TeV, normalising the Higgs cross section to 54.67 pb [11] and using branching ratios to $ee\mu\mu$ and $4e/4\mu$ of $5.9\times 10^{-5}$ and $6.5\times 10^{-5}$, respectively [12]. Hadronisation and parton showering are performed with Pythia [13] and detector simulation with Delphes [14] using the default card for the HL-LHC. Electrons are selected in the pseudorapidity region $|\eta|\leq 2.5$, and with transverse momentum $p_{T}\geq 7$ GeV. Muons are selected with $|\eta|\leq 2.7$ and $p_{T}\geq 5$ GeV. We require two opposite-sign same-flavour leptons. The three higher-$p_{T}$ leptons must satisfy minimum thresholds of 20, 15 and 10 GeV, respectively. These event selection requirements are quite close to those in Ref. [15]. In addition, we select a window $120\leq m_{4\ell}\leq 130$ GeV for the four-lepton invariant mass. In this region, the background has a size around 1/4 of the signal [15, 16]. The expected number of events in the $ee\mu\mu$, $4e$ and $4\mu$ final states are 1144, 379 and 1106, respectively. At the reconstructed level, $c_{111-1}$ and $c_{1010}$ exhibit moderate deviations from the parton-level values. We collect them in Table 1. For completeness we also give the values for the electroweak four-lepton background, calculated at the tree level, in the mass window $120\leq m_{4\ell}\leq 130$ GeV.
@@
A final comment on the electroweak four-lepton background is in order. In the region of interest it is quite small, as mentioned; the coefficients are also quite similar in $ee\mu\mu$ and $4e/4\mu$ final states, as seen in Table 1. And they are of similar magnitude as the ones for $H\to ZZ$. Therefore, a background subtraction, which is out of the scope of the present work, seems feasible, and will only result in a mild increase of the statistical uncertainty in $c_{111-1}$ and $c_{1010}$, and a corresponding decrease of the significance with respect to the results presented here. Next-to-leading order corrections for the $ee\mu\mu$ final state are known [19], and provide a small shift in $c_{111-1}$ and $c_{1010}$ which is below the statistical uncertainty.